### v1

# Genetic Instruct: Scaling up Synthetic Generation of Coding Instructions for Large Language Models

Abstract
Large Language Models (LLMs) rely on instruction samples for alignment, but creating these datasets poses challenges, particularly in expert-dependent tasks like coding, which can be cost-prohibitive. One approach to mitigate these challenges is synthesizing data using another LLM. In this paper, we introduce a scalable method for generating synthetic instructions to enhance the code generation capability of LLMs. The proposed algorithm, Genetic-Instruct, mimics evolutionary processes, utilizing self-instruction to create numerous synthetic samples from a limited number of seeds. Genetic-Instruct is designed for efficient scaling of the generation process. Fine-tuning multiple coding LLMs with the synthetic samples demonstrates a significant improvement in their code generation accuracy compared to the baselines.
spacing=nonfrench
Genetic Instruct: Scaling up Synthetic Generation of
Coding Instructions for Large Language Models
Somshubra Majumdar*, Vahid Noroozi thanks: Equal contribution, Sean Narenthiran, Aleksander Ficek, Jagadeesh Balam, Boris Ginsburg NVIDIA Santa Clara, CA 15213, USA {smajumdar,vnoroozi,snarenthiran}@nvidia.com
1 Introduction
Large Language Models (LLMs) have demonstrated significant advancements in programming and coding problems and can be readily used as code assistant models (Liang et al., 2024). To fully harness their capabilities, a process known as alignment (Ouyang et al., 2022) is necessary, which relies on paired instruction-solution examples to guide the model’s behavior and responses. One challenge in alignment is creating diverse and complex instructions, which can be costly, particularly in domains like coding that require specialized expertise. One strategy to address the scarcity of instructional examples for training is to generate synthetic instructions using another LLM. Research has shown that synthetic instructions can be effective not only for coding applications (Luo et al., 2024; Rozière et al., 2024; Yu et al., 2024) but also for general tasks (Wang et al., 2023; Honovich et al., 2023; Xu et al., 2024). This approach can be seen as a form of knowledge distillation or transfer from a more capable LLM to another.
Figure 1: Genetic-Instruct Process. A small seed population is operated on by an Instructor-LLM to perform crossover/mutation operations. A Coder-LLM then generates an appropriate code solution, which is judged for correctness by a Judge-LLM. Once a required population size is reached, the samples are de-duplicated and form the final population.
In this paper, we introduce a scalable algorithm named Genetic-Instruct, illustrated in Figure 1. Drawing inspiration from evolutionary algorithms, Genetic-Instruct is designed to generate synthetic instructions for coding problems. The algorithm begins with a small set of high-quality initial instruction samples and employs some LLMs to generate instructions and their corresponding code outputs through two evolutionary operations: crossover and mutation. The crossover operation is adapted from self-instruct (Wang et al., 2023), which prompts an LLM to generate new examples based on the context of a few-shot examples. The LLM generates new samples using seed instructions as a foundation, akin to a crossover operation. In the mutation operation, an LLM is given an instruction and asked to evolve it into another instruction based on predefined rules (Luo et al., 2024). Subsequently, another LLM proficient in coding generates the output or solution to the instruction, typically comprising coding solutions.
We also introduce a fitness function that employs an LLM to assess the correctness and quality of the generated instruction and its solution. Samples that pass these evaluations and checks are added to the population pool, and the evolutionary process continues until the target population size is reached. The entire pipeline is designed for efficient parallel execution with multiple colonies of populations. We extend previous work by providing a solution that enables existing operators with scalability as needed.
We generate a large dataset of synthetic coding instructions using the proposed Genetic-Instruct algorithm, starting from a limited set of coding problems. We then assessed the effectiveness of this dataset by training multiple open-source software (OSS) LLMs. Our work is dedicated to enhancing OSS models, and we strictly avoid using any closed-source data or models. The coding models trained with our synthetic dataset were evaluated across a range of coding benchmarks, showcasing the superior performance of our approach compared to other algorithms designed for generating synthetic coding instructions.
2 Previous Works
Synthetic data generation has emerged as a viable alternative to the time-intensive and expensive process of collecting data from human interactions with LLM systems. One notable method for automated data generation is Self-Instruct (Wang et al., 2023), which employs a pre-trained base LLM to generate instruction-output pairs from a small seed dataset. The resulting dataset is then used to fine-tune the base LLM, resulting in an Instruction-tuned LLM. This work is not focused on coding problems and generate instruction for general tasks.
A limitation of Self-Instruct is that the synthesized samples are often simple and lack diversity, failing to generate challenging instructions that require additional steps to arrive at the solution. To address this, Xu et al. (2024) introduced Evol-Instruct, an enhancement of Self-Instruct that integrates the concept of mutation of instructions to create more complex tasks. By introducing meta-instructions that increase the reasoning depth, add constraints to solutions, or evolve instructions with similar concepts (in-breadth evolution), Evol-Instruct produces a diverse set of complex instructions that significantly improve the performance of the instruction-tuned model. This concept was further adapted for the coding domain by Luo et al. (2024) in a variant called WizardCoder, where models trained on these datasets exhibit enhanced coding capabilities. Our work extends WizardCoder mutation by supporting mutating multiple generations of the same sample and contrarily open-sourcing our work.
While Self-Instruct and Evol-Instruct primarily focus on synthesizing code generation-style instruction-solution pairs, Yu et al. (2024) introduced the CodeOcean dataset, which adopts a novel approach to synthesizing complex tasks for code models. The models trained on CodeOcean, referred to as WaveCoder, are adept at code completion, documentation, fixing, and translation. Contrary to previous methods that construct instructions based on in-context seed instructions, the CodeOcean dataset leverages code snippets from existing codebases as context. The LLM is then tasked with constructing a valid instruction, one of the four tasks, followed by an appropriate solution, leading to a more diverse and challenging set of instructions. Other works such as PromptBreeder (Fernando et al., 2023) combine mutation and crossover but at the prompt level and then operating on the provided instruction to create a new solution. Our work mutates instructions by performing Wizard-Coder (and potentially Wave-Coder) on the given instruction as opposed to swapping a mutated prompt for a previous task prompt.
Holistically, our work does not replace previous methods but instead utilizes them as operators in a pipeline to automatically construct novel data with minimal human involvement and trivial amounts of seed data. This is due to features of massive parallel scaling, operator neutrality and compound mutations.
3 Genetic-Instruct
We introduce Genetic-Instruct, an algorithm inspired by the population-based evolutionary approach of genetic algorithms. This algorithm employs two primary evolutionary operations of mutation and crossover to evolve and generate new generations from an initial population. The initial population, termed Generation 0, comprises a limited set of high-quality seed instructions. These seed instructions undergo a series of evolutionary operations, mainly mutation, crossover and selection, to transform them into new instructions. In this study, the operations of mutation, crossover, and selection are executed by leveraging large language models (LLMs) and enhancing their output with in-context learning (Brown et al., 2020).
We designate the LLM responsible for instruction generation as the Instructor-LLM. Upon generating a new instruction, another LLM, referred to as the Coder-LLM, is tasked with producing the code corresponding to this new instruction. The Instructor-LLM and Coder-LLM can either be distinct models or the same model but with different prompts. The newly generated instruction and its associated code constitute a new coding instruction, which can be utilized for training. However, there may be instances where the generated code does not fully address the provided question, or the question itself may be poorly formulated. To assess the quality of the new coding instruction, we employ another LLM, termed the Judge-LLM, to evaluate the correctness of the instruction and its code. If a sample passes this quality assessment, it is added to the pool of instructions and may be selected as the seed instruction for the next batch of synthesized samples.
The entire process is iterated multiple times to synthesize samples until the desired population size is achieved. Subsequently, a deduplication process is applied to minimize duplicate instructions in the current dataset. This resulting population is then labeled as a generation, and the entire pipeline can be repeated by considering this generation as the initial population for the next generation. The complete pipeline is illustrated in Figure 1, and the procedure for generating one generation is detailed in Algorithm 9.
3.1 Selection Process
Initially, from the instruction set of the initial population ($S$) (also called the seed dataset), we randomly select a batch of instructions ($S^{\prime}$) with replacement. This batch is then subjected to either mutation or crossover steps as detailed below, with the generations denoted by $S^{\prime}_{gen}$. In subsequent iterations, we augment the initial population with the generations from the previous iteration, such that $S:=S\cup S^{\prime}_{gen}$. This process continues until a sufficient number of samples have been synthesized for the current generation.
3.2 Mutation Operation
The mutation operation is inspired by an adaptation of the Evol-Instruct algorithm, as devised by Xu et al. (2024), and further extended by WizardCoder (Luo et al., 2024) to facilitate instruction generation for code models. Evol-Instruct evolves an instruction into another using an LLM based on predefined tasks. For a sample selected for mutation, we randomly choose one of the five tasks defined and apply the mutation to generate a new instruction. We employ the same five tasks introduced by Luo et al. (2024), with minor prompt modifications to suit our Instructor-LLM. Details on the mutation prompts are provided in Figure 2.
3.3 Crossover Operation
The crossover operation in Genetic-Instruct is influenced by concepts introduced in Self-Instruct (Wang et al., 2023) and Unnatural Instructions (Honovich et al., 2023). It combines multiple genes (instructions) and employs the Instructor-LLM to generate new populations from the provided few-shot example instructions. To enhance the efficiency of the crossover operation, we provide multiple seed instructions and request the model to generate multiple diverse new instructions based on the provided examples in a single Instructor-LLM call. Simple post-processing, such as rejecting all samples that don’t pass the Abstract Syntax Tree checks, is then applied to filter out incorrect instructions. The prompt for the crossover operation is depicted in Figure 3.
3.4 Code Generation
After the Instructor-LLM generates a batch of new instructions, they are passed to the Coder-LLM to generate the corresponding code solutions. The Coder-LLM should be proficient in coding tasks to ensure the generation of high-quality solutions. However, some generated code may not be parseable or compilable. Therefore, we filter out solutions whose code segments cannot be parsed by the corresponding language’s parser/compiler. While determining the correctness of code by execution is the ideal case, it is challenging due to various factors, such as language constraints, missing dependencies, or having to integrate the current solution into a much larger codebase that may not be available in its entirety. The prompt used in this step is illustrated in Figure 4.
3.5 Fitness Function
In genetic algorithms, once candidate solutions are produced, they are scored using a fitness function in order to discard candidates that score lower compared to others. Similarly, we employ a Judge-LLM to assign a binary score indicating whether a candidate code solution meets the requirements of the corresponding instruction. The Judge-LLM is provided with an instruction and its code solution to determine the correctness of the instruction and its corresponding solution. To enhance the Judge-LLM’s performance, we employ techniques such as in-context learning with few-shot examples and Chain-of-Thought (Wei et al., 2022) prompting to facilitate reasoning before making a binary decision. The prompt for the Judge-LLM is depicted in Figure 5.
3.6 Scaling Up the Process
An advantage of genetic algorithms is their inherent capacity for parallelization. When utilizing computationally intensive LLMs for sample generation, it is crucial to exploit this parallel architecture. We efficiently compute multiple colonies of populations in parallel processes or compute nodes and synchronize them periodically. These colonies are evolved and populated independently, starting from the same seed population. Upon reaching the desired size, the colonies are merged into a single population, where deduplication is performed to obtain a unified dataset, termed one generation. We use only the instruction itself as the deduplication criterion to eliminate redundancies across different colonies. Deduplication is performed using the MinHash + LSH algorithm, implemented in text-dedup library. This generation serves as the shared seed data for the next generation, a process we refer to as inter-node parallelization.
Within each colony, using large batch sizes and selecting a single task for each algorithm iteration leads to redundancy in the generated instructions, and increases duplication. To address this, we employ sub-processes to compute multiple sub-colonies within each colony, each possessing a small batch size but capable of selecting diverse tasks. This approach maintains similar computational and memory requirements as a single large batch of single-task samples while enhancing the diversity of generated solutions, thereby reducing duplication. This process is termed intra-node parallelization.
4 Experiments
In all experiments, the models are assessed using four benchmark datasets: HumanEval (Chen et al., 2021), MBPP (Odena et al., 2021), HumanEval+, and MBPP+ (Liu et al., 2023). The MBPP+ and HumanEval+ datasets from the EvalPlus benchmark are extensions of the MBPP and HumanEval test sets, respectively. These extensions are enriched with additional test cases to ensure the correctness and accuracy of the generated code. All coding evaluations are conducted using greedy decoding. In all reported results, an asterisk (*) indicates that we computed the results ourselves, as the original works did not report these numbers.
We use the Tiger-Leetcode collection [^1] TigerResearch/tigerbot-kaggle-leetcodesolutions-en-2k as the initial population in most of the experiments. This collection serves as the seed dataset for the generation process and contains interview-style coding questions and solutions.
In all the experiments, we employed the same generation model as Instructor-LLM, Coder-LLM, and Judge-LLM. Since we are evaluating our approach solely on Python coding benchmarks, we explicitly limit our generated solutions to Python by requesting the models to generate only Python code. After the code is generated by Coder-LLM, we verify its syntactical correctness using Python’s ast package, regardless of its executability.
4.1 Experimental Settings
Table 1: Comparison of models fine-tuned with Genetic-Instruct with other baselines. Cells marked with $\times$ indicate result could not be calculated since the model is not available. The column labelled as OpenAI indicates whether the synthetic data is generated with any of the OpenAI’s models like ChatGPT. Best results for each base model are bolded. Cells marked with ✝ are evaluated on the original 500 un-sanitized MBPP dataset and aren’t directly comparable to the MBPP’s result calculated by EvalPlus.
Size Base Model Generation Algorithm OpenAI HumanEval HumanEval+ MBPP MBPP+
47B Mixtral-8x7B - N 40.2 34.3 60.7 50.1
- GPT-4-Turbo - Y 85.4 81.7 83.5 70.7
7B DeepSeek-Base WaveCoder Y 64.0 $\times$ 62.8 ✝ $\times$
7B CodeLlama-Python WaveCoder Y 48.1 $\times$ 47.2 $\times$
7B CodeLlama-Python WizardCoder Y 55.5 45.9 59.9 49.1
34B CodeLlama-Python WizardCoder Y 73.2 64.6 73.2 59.9
7B DeepSeek-Base - N 47.4 39.5 60.6 52.1
Genetic-Instruct N 57.9 51.8 70.4 59.1
[…]
CodeLlama-Instruct N 34.8 30.2 * 44.4 38.7 *
Genetic-Instruct N 50.6 47.6 65.2 55.9
34B CodeLlama-Python - N 53.7 46.2 * 56.2 49.3 *
CodeLlama-Instruct N 41.5 33.7 * 57.0 50.5 *
Genetic-Instruct N 59.1 53.7 67.7 57.1
70B CodeLlama-Python - N 57.3 51.3 * 65.6 53.4 *
CodeLlama-Instruct N 67.8 58.3 * 62.2 51.8 *
Genetic-Instruct N 57.9 55.5 69.2 58.6
We trained the base models using the AdamW optimizer (Kingma and Ba, 2015) with a learning rate of 5e-6 and zero weight decay for two epochs. The learning rate scheduler employed is CosineAnnealing with 100 warmup steps. The final model is obtained by averaging the last four checkpoints, which are saved every 500 steps. All models are trained using tensor parallelism, and BF16 precision is utilized to accelerate the training process.
For large throughput inference with large effective batch sizes, we employ vLLM (Kwon et al., 2023) as our inference engine. Nucleus sampling (Holtzman et al., 2020) is used for decoding, with a temperature of 0.95 for the Instructor-LLM and a temperature of 1.0 for the Coder-LLM and Judge-LLM. To enhance GPU utilization and expedite the generation process, we implement both inter-node and intra-node parallelizations. Specifically, we use 20 colonies for each generation in inter-node parallelization and 10 sub-colonies per colony in intra-node parallelization. A maximum sequence length of 1024 is set for all LLMs to optimize the generation speed and memory allocation.
For the mutation operation, a batch size of 100 per sub-colony is employed. In the crossover operation, a batch size of 10 per sub-colony is used, as we find that the model approximately generates 10 unique instructions on average per sample, resulting in a similar number of generated samples per batch. We randomly choose to perform either crossover or mutation on each batch. The Instructor-LLM uses 3-shot samples for in-context learning and is instructed to generate up to 20 new instructions.
The deduplication process is conducted using the MinHash algorithm from the TextDedup library [^2] https://github.com/ChenghaoMou/text-dedup with a 5-gram overlap, 250 permutations, and a threshold of 0.7. The synthetic dataset generated by our proposed algorithm, Genetic-Instruct, comprises approximately 1.3 million coding instructions, which are divided into different subsets for evaluation. More details on the generated dataset can be found in Appendices F and G.
Table 2: Comparing the effectiveness of different operations in the Genetic-Instruct algorithm. Model CodeLlama-Python 7B has been used as the base model in these experiments. Bolded cells indicate best score for a given dataset size.
Population Size Generation Algorithm HumanEval HumanEval+ MBPP MBPP+
- CodeLlama-Python 7B 38.4 35.2 * 47.6 40.5 *
Cross-Over Only 47.0 41.5 55.4 44.4
2K Mutation Only 40.2 34.8 60.4 47.1
Genetic-Instruct 46.3 41.5 56.4 45.4
Cross-Over Only 48.2 45.7 59.1 51.6
75K Mutation Only 43.9 39.0 63.2 51.9
Genetic-Instruct 49.4 45.7 62.7 52.9
Cross-Over Only 51.7 47.6 62.2 52.9
250K Mutation Only 47.6 42.1 62.7 52.1
Genetic-Instruct 51.8 47.6 63.9 54.6
Table 3: Ablation study on the effect of the generator model on data generation quality. We did the evaluations on two base models: DeepSeek Coder 6.7B and 33B. The first three rows are the generation models used for data synthesis.
Model Dataset Size HumanEval HumanEval+ MBPP MBPP+
Mixtral 8x7B - 40.2 34.3 60.7 50.1
Mixtral 8x22B - 75.3 71.1 71.9 63.6
Llama 3 70B - 77.1 71.2 80.3 67.8
DeepSeek Coder 6.7B Base 47.4 39.5 60.6 52.1
Mixtral 8x7B 900K 54.9 51.8 71.4 60.4
Mixtral 8x22B 377K 66.5 62.8 72.2 60.4
Mixtral 8x22B 873K 68.3 66.5 69.9 59.9
Mixtral 8x22B 1.3M 70.7 67.7 75.4 63.7
Llama 3 8B 400K 63.4 57.9 70.2 58.6
Llama 3 8B 1.2M 64.6 61.0 69.9 57.1
Llama 3 70B 767K 66.5 62.8 72.2 61.4
Llama 3 70B 1.8M 70.7 66.5 75.4 63.2
DeepSeek Coder 33B Base 51.2 44.5 65.4 56.8
Llama 3 70B 767K 70.1 65.9 77.9 64.7
Llama 3 70B 1.8M 73.8 68.9 77.7 64.7
4.2 Performance Evaluation
In this section, we evaluate the effectiveness of our proposed approach to generate synthetic SFT samples that can enhance the coding capabilities of LLMs. We used CodeLLama-Python-7B (Rozière et al., 2024) and DeepSeek-Coder-6.7B (Guo et al., 2024) as the base models and employed Mixtral 8X7B (Jiang et al., 2024) as the Instructor-LLM, Coder-LLM, and Judge-LLM. The results, along with the performance of other coding models, are reported in Table 1. For cases where no generation algorithm is used, the results of the base models are reported. All results are obtained using greedy decoding with the pass@1 metric. The base models are fine-tuned with a subset of 405K samples from all Genetic-Instruct’s generations, while the WizardCoder and WaveCoder models are trained with datasets of 78K and 20K samples, respectively.
The results show that training models with Genetic-Instruct significantly improves performance on all benchmarks compared to the base models. This demonstrates the effectiveness of the Genetic-Instruct approach in enhancing the code generation capabilities of LLMs. The results for different sizes of the CodeLlama-Python model indicate that as the size of the base model increases, stronger performances can be achieved on most benchmarks. It should also be noted that HumanEval+ and MBPP+ are more trustworthy benchmarks as they are more robust to incorrect codes.
The results of GPT-4-Turbo and Mixtral 8x7B are also reported as references to illustrate the performance gap between OpenAI’s model and Mixtral. We used the Mixtral 8X7B as the generation model, which is a permissive model, whereas other approaches like WaveCoder or WizardCoder have used OpenAI’s models, which are not viable for the open-source community. To ensure a fair comparison between Genetic-Instruct and other synthetic data generation methods, we conducted more controlled experiments in Section 4.3.
It should be noted that any involvement of OpenAI’s models (Brown et al., 2020) in data synthesis has a marked positive effect on the downstream model, enabling substantial improvements in all models via data distillation. It can be seen that for WaveCoder and WizardCoder, results on all benchmarks are significantly improved compared to their base models. However, Genetic-Instruct surpasses these 7B models on EvalPlus datasets, indicating that such data distillation is not particularly robust. These results further suggest that while the choice of generator LLM is crucial in improving the results of downstream models, larger datasets can also provide more robust results.
An interesting observation of CodeLlama-70B’s results is its ability to surpass the Mixtral-8x7B model scores on all benchmarks despite using the relatively weaker Mixtral-8x7B models to generate data. The CodeLlama-70B-Python model surpasses Mixtral-8x7B’s scores on all benchmarks and shows further improvements after training on data generated by Mixtral-8x7B. This indicates that relatively weaker models can generate sufficiently diverse and challenging data, which, when applied to stronger domain-specific models, can outperform both. Essentially, we observe that improvements transfer from a relatively weaker model to a stronger model. This goes against the general trend of data distillation, which is usually from a stronger model to a weaker model. This result may enable an iterative self-improvement loop, such that the model fine-tuned on Genetic-Instruct data can be used to generate another dataset with higher quality for subsequent generations of models, similar to Noisy Student Training proposed by Xie et al. (2020).
4.3 Ablation Study
In this ablation study, we evaluate the effectiveness of the operations used in Genetic-Instruct on the performance of the generated samples. We perform three data generation ablations, labeled Crossover Only, Mutation Only, and Genetic-Instruct, which represent data generated using the Genetic-Instruct algorithm but with one task disabled (either crossover or mutation, respectively). We generate 250K samples for each dataset, and then uniformly sample without replacement subsets of 2K and 75K from this total dataset. We then train a CodeLlama-Python-7B model on each dataset and evaluate their performance. This ablation provides an objective comparison of Genetic-Instruct generation with alternatives Self-Instruct (Wang et al., 2023) and WizardCoder (Luo et al., 2024), which use crossover only and mutation only, respectively. By controlling variables such as compute, amount of data, generation model, and base model, we demonstrate a head-to-head comparison of generation techniques.
As seen in Table 2, crossover provides substantial improvements over the base model compared to mutation alone; however, both are necessary to achieve the best results. At just 2K samples, there is no clear winner between the three approaches, potentially due to the noisy sampling of 2K samples out of 250K, but all approaches improve on the baseline. As the dataset size increases to 75K and 250K, the combination of both crossover and mutation surpasses either separately, and this trend continues at larger data scales as measured on the EvalPlus benchmark.
Table 4: The effect of the population size on the performance of the Genetic-Instruct.
Dataset Size HumanEval HumanEval+ MBPP MBPP+
- 38.4 35.2 * 47.6 40.5 *
2K 46.3 41.5 56.4 45.4
75K 49.4 45.7 62.7 52.9
250K 51.8 47.6 63.9 54.6
405K 50.6 47.6 65.2 55.9
900K 51.2 48.2 64.9 56.4
4.4 Influence of the Generator Model on the Performance
A critical component of the Genetic-Instruct algorithm is the choice of the Instructor-LLM, Coder-LLM, and Judge-LLM. A powerful Instructor-LLM can generate more diverse instructions, a stronger Coder-LLM can produce code that is more likely to be correct and free from syntactic errors, and an effective Judge-LLM can better discern the quality of samples.
In Table 3, we utilized three different models (Mixtral 8x7B, Mixtral 8x22B, and Llama-3 70B) to generate a large amount of data in two to three generations and fine-tuned two base DeepSeek Coder models on them for comparison. We found that the quality of the final model strongly correlates with the quality of the data generated by it, and that different dataset sizes can achieve similar scores when two sufficiently powerful models are compared. An important observation is that while Mixtral 8x22B data shows strong results, obtaining similar scores to Llama 3 70B generated data in fewer samples, it required three generations of data to achieve similar scores to two generations for Llama 3 70B, thereby requiring significant additional compute.
Furthermore, when scaling up the dataset, the gap between the original and EvalPlus benchmarks is substantially reduced, as seen in the CodeLlama-7B experiments. Similarly, the results on CodeLlama-70B show a marginal improvement on the original HumanEval dataset but a substantial improvement on the EvalPlus counterpart. This indicates that despite saturation on the original benchmarks, training on more synthetic data improves the robustness of the models on extended benchmarks.
Table 5: Performance of the dataset generated by Genetic-Instruct with different seed datasets. CodeLlama-Python 7B has been used as the base model in the experiments.
Seed Dataset Size HumanEval HumanEval+ MBPP MBPP+
No SFT - 38.4 35.2 * 47.6 40.5 *
Code Alpaca 20K 49.4 45.7 63.4 53.4
Tiger-LeetCode 2K 50.6 47.6 65.2 55.9
4.5 Effectiveness of Scaling up the Population
In this section, we examine the impact of the size of generated samples on model performance. We generated 250K samples with Mixtral 8x7b and then uniformly sampled without replacement subsets of 2K and 75K from it. Subsequently, we trained a CodeLlama-Python 7B model on each dataset and evaluated their performance. The results are shown in Table 4. As the dataset size increases, performance generally improves, indicating that scaling synthetic data generation with our proposed algorithm can be beneficial. This trend is more evident in the MBPP+ and HumanEval+ datasets, showing that more synthetic data generated by Genetic-Instruct can enhance the quality and correctness of the codes the model generates.
4.6 Effect of the Initial Population
In this section, we evaluate the Genetic-Instruct algorithm on different datasets as the initial population to demonstrate the effectiveness of the proposed approach. We utilize CodeLlama-7B and two seed datasets: Tiger-leetcode and Code Alpaca333sahil2801/CodeAlpaca-20k. We generated roughly 400K samples with Genetic-Instruct using Code Alpaca as the seed set for comparison against our Tiger-leetcode baseline. Code Alpaca is an instruction dataset of around 20K samples created by ChatGPT 3.5. Table 5 shows the results of different seed datasets when used with Genetic-Instruct.
We observe that our method is robust to the seed dataset, capable of generating instruction data of similar quality for both seed sets, and significantly surpasses the base model’s results. There is a modest gap between the two seed datasets, which we attribute to the simplicity of instructions in Code Alpaca. It is important to note that Genetic-Instruct does not utilize the code solutions of the seed dataset when creating new instructions via the Instructor-LLM. We conjecture that the relative difficulty and complexity of Tiger-Leetcode’s questions lead to a more diverse synthetic dataset.
Given the similar performance observed between the Code Alpaca seeds generated by ChatGPT-3.5 and Tiger-leetcode, it appears that the success of the approaches utilizing OpenAI’s models is primarily due to the model used for generating instructions and codes, rather than the seeds themselves. This observation underscores the importance of the quality of LLMs employed in the generation process.
5 Conclusion
In this paper, we introduced Genetic-Instruct, a novel algorithm inspired by evolutionary principles, designed to generate large-scale synthetic coding instructions for large language models (LLMs). The algorithm utilizes two evolutionary operations, crossover and mutation, applied iteratively to a population pool to produce new instruction samples until the desired size is reached. Genetic-Instruct is specifically engineered to support parallel generation, making it a scalable solution for synthetic data creation. We benchmarked our approach against several baseline methods, demonstrating that Genetic-Instruct can generate instructions that lead to higher accuracy on code generation benchmarks. We also showed that increasing the size of the generated samples improves the performance of the final models. Furthermore, we evaluated the algorithm’s performance across various initial populations and base models, showcasing its robustness with respect to seed instructions and base LLMs.
Appendix A Mutation Prompts
Mutation Prompt Please increase the difficulty of the given programming test question a bit. Do not provide any hints, solutions or outputs. Only one new instruction is allowed. You can increase the difficulty using, but not limited to, the following methods: {method} Original Instruction: {instruction} New Instruction:
Operation: Constraint Rewrite the original instruction, adding new constraints and requirements, with approximately 10 additional words.
Operation: Deepening Write the original instruction. Then, replace a commonly used requirement in the programming task with a less common and more specific.
Operation: Erroneous Code Write the original instruction. Then provide a piece of wrong python code as a reference solution to increase misdirection. Your wrong reference solution should start with "Reference Solution (Wrong)", marked in “‘ blocks. Finally, ask to write the correct solution for the instruction. Do NOT provide the correct solution.
Operation: Reasoning Write the original instruction after the new instruction. Then, if the original instruction can be solved with only a few logical steps, please add more reasoning steps after the original instruction. Do NOT provide any reason or explanation.
Operation: Task Complexity Write the original instruction after the new instruction. Then propose higher time or space complexity requirements, but please refrain from doing so frequently.
Figure 2: Prompt template for mutation operation
Appendix B Crossover Prompt
Cross-over Prompt You are asked to come up with a set of 20 diverse code generation task instructions. These task instructions will be given to a GPT model and we will evaluate the GPT model for completing the instructions. Here are the requirements: 1. Try not to repeat the verb for each instruction to maximize diversity. 2. The language used for the instruction also should be diverse. For example, you should combine questions with imperative instructions. 3. The type of instructions should be diverse. The list should include diverse types of programming tasks like open-ended generation, classification, editing, optimization etc. 4. A GPT language model should be able to complete the instruction. 5. The instructions should be in English. 6. The instructions should at least 1 to 2 sentences long. Either an imperative sentence or a question is permitted. 7. You should generate an appropriate input to the instruction. The input field should contain a specific example provided for the instruction. It should involve realistic data and should not contain simple placeholders. The input should provide substantial content to make the instruction challenging but should ideally not exceed 100 words. 8. Not all instructions require input. For example, when a instruction asks about some general information, "write a program to load a file.", it is not necessary to provide a specific context. In this case, we simply put "⟨noinput⟩" in the input field. 9. The output should be an appropriate response to the instruction and the input. 10. All tasks should be coding or programming-related. List of 20 tasks:
Few-Shot Examples ### 1. Instruction: Convert a Binary Search Tree to a sorted Circular Doubly-Linked List in place. You can think of the left and right pointers as synonymous to the predecessor and successor pointers in a doubly-linked list. For a circular doubly linked list, the predecessor of the first element is the last element, and the successor of the last element is the first element. We want to do the transformation in place. After the transformation, the left pointer of the tree node should point to its predecessor, and the right pointer should point to its successor. You should return the pointer to the smallest element of the linked list. 1. Input: root = 4,2,5,1,3 ### 2. Instruction: $\cdots$ $\cdots$ ### 4. Instruction:
Figure 3: Prompt template for the cross-over operation with few-Shot in-context learning
Appendix C Prompt Used for Generating Code with Coder-LLM
Python Code Generation Prompt You are an expert in Python coding. Using only Python code, write the correct solution that answers the given coding problem. {instruction} Answer:
Figure 4: Prompt template for code Generation with Coder-LLM
Appendix D Fitness Prompt for Judge-LLM
[…]
Figure 5: Prompt template for code quality judgement with Judge-LLM
Appendix E Evaluation Prompt
Evaluation Prompt Template Below is an instruction that describes a task. Write a response that appropriately completes the request. ### Instruction: {input} ### Response:
Figure 6: Prompt template for code evaluation on HumanEval and MBPP
Appendix F Distribution of the Tasks
Figure 7: Task distribution of the generated instructions on roughly 600K samples of Genetic-Instruct out of 900K. Considering the domain of seed data, we select keywords that correspond to coding challenges. It is not an exhaustive list of the tasks and they can be overlapping.
Appendix G Length Distribution
Figure 8: Length distribution of the generated instructions, and code outputs on 900K samples of Genetic-Instruct. Instruction lengths are clipped to 300 words, while code outputs are clipped at 500 words for better visualization.
Appendix H Genetic-Instruct Algorithm
Genetic-Instruct Algorithm {python} import numpy as np task_ids = 0: "cross-over", 1: "mutation" task_prob = [0.5, 0.5] # 50 NUM_SAMPLES, NUM_COLONIES, NUM_SUBCOLONY = 500_000, 20, 10 def genetic_instruct(seed_instructions: List[Dict[str, str]]): """ Genetic Instruct for one generation, run on multiple nodes """ max_samples = NUM_SAMPLES / NUM_COLONIES current_samples = [] all_samples = [v for v in seed_instructions] while len(current_samples) <= max_samples: # Parallelize across NUM_SUBCOLONY processes (sub-colonies) futures = parallelize(genetic_instruct_one_iter, (all_samples,) for _in range(NUM_SUBCOLONY) )) results = await_all(futures) # wait for all sub-colonies to finish for result in results: current_samples.extend(result) all_samples.extend(result) # Deduplicate across all colonies current_samples = global_deduplicate_across_nodes(current_samples) write_dataset(current_samples) def genetic_instruct_one_iter(seed_instructions: List[Dict[str, str]]): """ One iteration of the Genetic Instruct algorithm Inputs: seed_instructions: List of (instruction, code) pairs from seed data + previous generations. Outputs: List of (instruction, code) pairs. """ # (1) Task Selection task_id = np.random.choice(task_ids.keys(), p=task_prob) # (2) Instruction Generation - sample a batch of seed data prompts = encode_inst_generation_prompts(seed_instructions, task_id) new_instructions = llm.generate(prompts) # (3) Code Generation - generate code for new inst, filter generations code_prompts = encode_codegen_prompts(new_instructions, task_id) code_generations = llm.generate(code_prompts) new_inst_code = filter_code_generations(new_instructions, code_generations) # (4) Code Quality Judgement - judge inst-code pairs for correctness judge_prompts = encode_judge_prompts(new_inst_code) code_judgements = llm.generate(judge_prompts) final_samples = [new_inst for(new_inst, judgement) in zip(new_instructions, code_judgements) if judgement == ’yes’] return final_samples
Figure 9: Algorithm of genetic-instruct as pseudo-code